# Impact of new measurements of light quarks at hadron colliders

### v1

@@ II Review of the new Drell–Yan measurements
For the neutral current Drell–Yan measurement, a recent method proposed in Ref. AFBFactorization enables a refined flavor decomposition through the forward-backward asymmetry ($A_{FB}$) spectrum in the $Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ Drell–Yan process. The $u$ ($\bar{u}$) and $d$ ($\bar{d}$) quark information inside $A_{FB}$ observable can be factorized into a well-defined structure parameter, $R$, which serves as a novel experimental observable and reflects unique information of the relative difference between $u$ and $d$ quarks.
@@
Figure 1: Comparison between experimental measurements and theoretical predictions obtained using the modern PDF sets CT18NNLO CT18NNLO , MSHT20 MSHT20 , and NNPDF4.0 NNPDF4 , for observables sensitive to the flavor composition of the proton in hadron collider environments. The left panel shows the D0 measurement of the inverse $1/R$ in the Drell–Yan process at $\sqrt{s}=1.96$ TeV, while the middle panel presents the CMS measurement of $R$ at $\sqrt{s}=8$ TeV in a similar process. The right panel shows the ATLAS measurement of the $W$ boson charge asymmetry $A^{W}_{\text{ch}}$ at $\sqrt{s}$ = 13 TeV.
@@ III Dataset PDF Correlation Patterns and Implications
Figure 2: $C_{H}(E,f_{\text{parton}})$ (parton = $u_{v}$, $d_{v}$, $u$, $d$, $\bar{u}$, and $\bar{d}$, respectively) as a function of $x$, showing the correlation between each dataset and the parton PDF $f_{\text{parton}}$.
@@
To identify which parton distribution functions are primarily responsible for the potential deviations observed among the three datasets, and to assess the consistency of their $x$-dependent correlation patterns, the $C_{H}(E,f)$ between the new datasets and the $u_{v}$, $d_{v}$, $u$, $d$, $\bar{u}$, and $\bar{d}$ PDFs are shown in Figure 2. The used quark PDFs are from CT18NNLO. The D0 1.96 TeV $R$ measurement (the blue curve), the CMS 8 TeV $R$ measurement (the green curve) and the ATLAS 13 TeV $A^{W}_{\text{ch}}$ measurement (the pink curve) are plotted together.
@@
Very similar patterns of behavior are observed among the different $C_{H}(E,f)$ distributions in Figure 2, reflecting the common physical sensitivity of the three new measurements. For the valence $u$ quark PDF ($u_{v}$, top-left panel), all three datasets yield positive $C_{H}(E,f)$ values in the $x$ region around 0.1, corresponding to the larger-$x$ parton in their kinematics. From the definition of $C_{H}$, a positive correlation generally indicates that minimizing $\mathcal{X}^{2}_{E}$ for a given dataset is accompanied by a reduction in the value of $f_{u_{v}}$ in this region. This consistent trend implies that all three measurements favor a lower $u_{v}$ compared with the CT18NNLO prediction at $x\sim 0.1$, in line with the observed excess of $R$ and the deficit in $A^{W}_{\text{ch}}$ discussed in Section II. A complementary pattern is seen for the valence $d$ quark PDF ($d_{v}$, top-right panel), where all datasets show negative $C_{H}(E,f)$ values in the same $x$ region. This generally indicates that these datasets tend to enhance $d_{v}$ in this region, together with a reduction in $u_{v}$, implying that the deviations in $R$ and $A^{W}_{\text{ch}}$ can be jointly interpreted as a preference for a larger $d_{v}/u_{v}$ ratio than modern PDF predictions.
@@
The middle-row panels show the correlations with the $u$ and $d$ quark PDFs. At $x\sim 0.1$, $u$ exhibits a positive correlation for all datasets, again indicating a preference for a downward shift in this range, while $d$ exhibits the opposite sign, indicating an enhanced $d$ quark density. These patterns persist for the antiquark PDFs ($\bar{u}$ and $\bar{d}$, bottom panels), where $\bar{d}$ follows the same trend as $u$, and $\bar{u}$ follows the same trend as $d$. These results reinforce the conclusion that the observed deviations in $R$ and $A^{W}_{\text{ch}}$ can be simultaneously explained by a shift in the valence quark ratio $d_{v}/u_{v}$, which equals $(d-\bar{d})/(u-\bar{u})$.
@@
In summary, the strong correlation patterns for $u_{v}$ and $d_{v}$, together with the consistent behavior in the $u$, $d$, $\bar{u}$, and $\bar{d}$ PDFs, provide quantitative evidence that the three new datasets are pulling the current PDFs in the same direction of increasing $d_{v}/u_{v}$ ratio at $x\sim 0.1$. This mutual agreement supports the conclusion that the observed deviations are unlikely to be independent statistical fluctuations, but rather reflect a coherent deviation with the central predictions of current PDFs.
@@ IV Compatibility and Tension Analysis via Weighted PDF Updating
To further assess the compatibility between the three new hadron collider measurements and the datasets currently input to modern global PDF fits, as well as to investigate their impact on the PDF updating results, the error PDF updating method package (ePumpepump ; epump2 is employed to update the CT18NNLO CT18NNLO PDFs. ePump facilitates efficient updating of the best-fit PDF set and Hessian eigenvector pairs of PDF sets (i.e., error PDFs) in light of new data by retaining all theoretical assumptions of the original global fit, such as the choice of parametrization and number of parameters. In Ref. epump2 , the ePump framework is validated through detailed comparisons with full global analyses, and its potential is illustrated via selected phenomenological applications relevant to the LHC. In this work, the probed $x$ regions and the related energy scale $Q$ regions lie within the validated applicability range of Ref. epump2 . Therefore, despite the fixed assumptions, the qualitative conclusions regarding the impact of the new measurements remain valid.
@@
To further illustrate the impact of the three new datasets on individual parton densities, we compare the original CT18NNLO PDF set with the updated version generated through ePump tool, where each dataset was assigned a weight of 15. Figure 4 shows the ratios of the updated PDFs to the original ones, together with their uncertainties, for six selected parton flavors: $u_{v}$, $d_{v}$, $u$, $d$, $\bar{u}$, and $\bar{d}$, all evaluated at $Q=100$ GeV.
The most pronounced changes appear in the valence quark PDFs within the $x\sim 0.1$ region, where all three datasets have their strongest sensitivity. Consistent with the $C_{H}(E,f)$ correlation patterns in Figure 2, the updated $u_{v}$ distribution exhibits a downward shift in this range, while $d_{v}$ shows a corresponding enhancement, jointly indicating a preference for a larger $d_{v}/u_{v}$ ratio. The same flavor-dependent tendency is also visible in the other quark PDFs: $u$ and $\bar{d}$ are reduced, whereas $d$ and $\bar{u}$ are enhanced, in agreement with the correlation analysis discussed in Section III. Outside the $x\sim 0.1$ region, the observed shifts are largely driven by the valence quark sum rules $\int_{0}^{1}u_{v}(x)\text{d}x=2$ and $\int_{0}^{1}d_{v}(x)\text{d}x=1$.

### v2

@@ II Review of the new Drell–Yan measurements
For the neutral current Drell–Yan measurement, a recent method proposed in Reference AFBFactorization enables a refined flavor decomposition through the forward-backward asymmetry ($A_{FB}$) spectrum in the $Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ Drell–Yan process. The $u$ ($\bar{u}$) and $d$ ($\bar{d}$) quark information inside $A_{FB}$ observable can be factorized into a well-defined structure parameter, $R$ CMS , which serves as a novel experimental observable and reflects unique information of the relative difference between $u$ and $d$ quarks.
@@
Figure 1: Comparison between experimental measurements and theoretical predictions obtained using the modern PDF sets CT18NNLO CT18NNLO , MSHT20 MSHT20 , and NNPDF4.0 NNPDF4 , for observables sensitive to the flavor composition of the proton in hadron collider environments. The left panel shows the D0 measurement of the inverse $1/R$ in the Drell–Yan process at $\sqrt{s}=1.96$ TeV, while the middle panel presents the CMS measurement of $R$ at $\sqrt{s}=8$ TeV in a similar process. The right panel shows the ATLAS measurement of the $W$ boson charge asymmetry $A^{W}_{\text{ch}}$ at $\sqrt{s}$ = 13 TeV. According to Equation 2, $R$ is defined as a function of $x_{1}$, with $x_{1}\sim 0.1$ being a valid approximation only in the rapidity ranges $1<|Y|<1.5$ for the D0 data and $1.25<|Y|<2.4$ for the CMS data. The complete set of rapidity bins in the left (middle) panel covers a wider range of $0.05<x_{1}<0.4$ ($0.005<x_{1}<0.1$) values.
@@ III Dataset PDF Correlation Patterns and Implications
Figure 2: $C_{H}(E,f_{\text{parton}})$ (parton = $u_{v}$, $d_{v}$, $d_{v}/u_{v}$, $u$, $d$, $d/u$, $\bar{u}$, $\bar{d}$ and $\bar{d}/\bar{u}$, respectively) as a function of $x$, showing the correlation between each dataset and the parton PDF $f_{\text{parton}}$ from CT18NNLO.
@@
To identify which parton distribution functions are primarily responsible for the potential deviations observed among the three datasets, and to assess the consistency of their $x$-dependent correlation patterns, the $C_{H}(E,f)$ between the new datasets and the $u_{v}$, $d_{v}$, $d_{v}/u_{v}$, $u$, $d$, $d/u$, $\bar{u}$, $\bar{d}$ and $\bar{d}/\bar{u}$ PDFs are shown in Figure 2. The used quark PDFs are from CT18NNLO. The D0 1.96 TeV $R$ measurement (the blue curve), the CMS 8 TeV $R$ measurement (the green curve) and the ATLAS 13 TeV $A^{W}_{\text{ch}}$ measurement (the pink curve) are plotted together. It should be emphasized that the exact definition of $R$, as given in Reference CMS , is employed in this chapter and in Chapter IV, rather than the approximate expressions presented in Equations 1 and 2.
@@
Very similar patterns of behavior are observed among the different $C_{H}(E,f)$ distributions in Figure 2, reflecting the common physical sensitivity of the three new measurements. For the valence $u$ quark PDF ($u_{v}$, top-left panel), all three datasets yield positive $C_{H}(E,f)$ values in the $x$ region around 0.1, corresponding to the larger-$x$ parton in their kinematics. From the definition of $C_{H}$, a positive correlation generally indicates that minimizing $\mathcal{X}^{2}_{E}$ for a given dataset is accompanied by a reduction in the value of $f_{u_{v}}$ in this region. This consistent trend implies that all three measurements favor a lower $u_{v}$ compared with the CT18NNLO prediction at $x\sim 0.1$, in line with the observed excess of $R$ and the deficit in $A^{W}_{\text{ch}}$ discussed in Section II. A complementary pattern is seen for the valence $d$ quark PDF ($d_{v}$, top-middle panel), where all datasets show negative $C_{H}(E,f)$ values in the same $x$ region. This generally indicates that these datasets tend to enhance $d_{v}$ in this region, together with a reduction in $u_{v}$, implying that the deviations in $R$ and $A^{W}_{\text{ch}}$ can be jointly interpreted as a preference for a larger $d_{v}/u_{v}$ ratio than modern PDF predictions, in agreement with the negative value of the valence $d$-to-$u$ ratio ($d_{v}/u_{v}$, top-right panel) at $x\sim 0.1$.
@@
The middle-row panels show the correlations with the $u$, $d$, and $d/u$ PDFs. At $x\sim 0.1$, $u$ exhibits a positive correlation for all datasets, again indicating a preference for a downward shift in this range, while $d$ exhibits the opposite sign, indicating an enhanced $d$ quark density. These patterns persist for the antiquark PDFs ($\bar{u}$, $\bar{d}$, and $\bar{d}/\bar{u}$, bottom panels), where $\bar{d}$ follows the same trend as $u$, and $\bar{u}$ follows the same trend as $d$ at $x\sim 0.1$. Consequently, these behaviors lead to a negative value of $d/u$ and a positive value of $\bar{d}/\bar{u}$ at $x\sim 0.1$, thereby reinforcing the conclusion that the observed deviations in $R$ and $A^{W}_{\text{ch}}$ can be simultaneously accounted for by a shift in the valence quark ratio $d_{v}/u_{v}$, defined as $(d-\bar{d})/(u-\bar{u})$.
@@
In summary, the strong correlation patterns for $u_{v}$, $d_{v}$, and $d_{v}/u_{v}$, together with the consistent behavior in the $u$, $d$, $d/u$, $\bar{u}$, $\bar{d}$, and $\bar{d}/\bar{u}$ PDFs, provide quantitative evidence that the three new datasets are pulling the current PDFs in the same direction of increasing $d_{v}/u_{v}$ ratio at $x\sim 0.1$. This mutual agreement supports the conclusion that the observed deviations are unlikely to be independent statistical fluctuations, but rather reflect a coherent deviation with the central predictions of current PDFs.
@@ IV Compatibility and Tension Analysis via Weighted PDF Updating
To further assess the compatibility between the three new hadron collider measurements and the datasets currently input to modern global PDF fits, as well as to investigate their impact on the PDF updating results, the error PDF updating method package (ePumpepump ; epump2 is employed to update the CT18NNLO CT18NNLO PDFs. ePump facilitates efficient updating of the best-fit PDF set and Hessian eigenvector pairs of PDF sets (i.e., error PDFs) in light of new data by retaining all theoretical assumptions of the original global fit, such as the choice of parametrization and number of parameters. In Reference epump2 , the ePump framework is validated through detailed comparisons with full global analyses, and its potential is illustrated via selected phenomenological applications relevant to the LHC. In this work, the probed $x$ regions and the related energy scale $Q$ regions lie within the validated applicability range of Reference epump2 . Therefore, despite the fixed assumptions, the qualitative conclusions regarding the impact of the new measurements remain valid.
@@
To further illustrate the impact of the three new datasets on individual parton densities, we compare the original CT18NNLO PDF set with the updated version generated through ePump tool, where each dataset was assigned a weight of 15. Figure 4 shows the ratios of the updated PDFs to the original ones, together with their uncertainties, for nine selected parton flavors: $u_{v}$, $d_{v}$, $d_{v}/u_{v}$, $u$, $d$, $d/u$, $\bar{u}$, $\bar{d}$ and $\bar{d}/\bar{u}$, all evaluated at $Q=100$ GeV.
The most pronounced changes appear in the valence quark PDFs within the $x\sim 0.1$ region, where all three datasets have their strongest sensitivity. Consistent with the $C_{H}(E,f)$ correlation patterns in Figure 2, the updated $u_{v}$ distribution exhibits a downward shift in this range, while $d_{v}$ shows a corresponding enhancement, jointly indicating a preference for a larger $d_{v}/u_{v}$ ratio. The same flavor-dependent tendency is also visible in the other quark PDFs: $u$, $\bar{d}$, and $\bar{d}/\bar{u}$ are reduced, whereas $d$, $\bar{u}$, and $d/u$ are enhanced, in agreement with the correlation analysis discussed in Section III. Outside the $x\sim 0.1$ region, the observed shifts are largely driven by the valence quark sum rules $\int_{0}^{1}u_{v}(x)\text{d}x=2$ and $\int_{0}^{1}d_{v}(x)\text{d}x=1$.